OpenStudy (anonymous):

determine whether the equation (16x^2-y^2=0) represents y as a function of x

OpenStudy (jim766):

try solving it for y

OpenStudy (jim766):

16x^2 - y^2 = 0 16x^2 = y^2 , now take the square root of both sides

OpenStudy (jim766):

take the square root of both sides

OpenStudy (anonymous):

I still don't get it

OpenStudy (jim766):

after taking the square root of both sides, you end up with 4x = y , do you know what kind of graph this would make?

OpenStudy (jim766):

do you have a graph calc?

OpenStudy (jim766):

y = 4x is a linear function....

OpenStudy (anonymous):

does this make it not a function of x then?

OpenStudy (jim766):

yep